# April 3, 2025

### v1

@@ Abstract.
In this paper, we establish the existence of solutions for time-dependent linearly viscoelastic bodies confined to a specified half-space. The confinement condition under consideration is of Signorini type, and is given over the boundary of the linearly viscoelastic body under consideration. In order to derive a suitable notion of solution that possesses a sufficient degree of regularity, we introduce the concept of “admissible” applied body force. In agreement with expectations, the variational problem that solutions must satisfy is a system of hyperbolic variational inequalities, and the displacement acceleration is a vector-valued measure in the sense of Dinculeanu.
@@ 1. Introduction
The contact of elastic and viscoelastic materials with rigid obstacles arises in many applicative fields, including biomedical engineering, aerospace design, and materials science. For instance, the motion of aortic heart valves - modelled as linearly elastic shells constrained within a spatial domain to avoid collisions - has inspired mathematical frameworks for obstacle problems in elasticity [22, 26, 39]. Similarly, the confinement of deformable structures like polymer films or viscoelastic membranes in prescribed half-spaces is critical for applications ranging from soft robotics to biomechanical implants [38].
@@
The mechanical deformation of linearly viscoelastic bodies is classically modelled using energy functionals from linearized elasticity theory [7, 22]. Recent advances extend these principles to dynamic and time-dependent settings. For example, Bock and Jarušek analysed unilateral dynamic contact problems for viscoelastic Reissner-Mindlin plates [2], while [3] established existence results for thermoelastic von Kármán plates vibrating against rigid obstacles. A key challenge in such problems lies in reconciling hyperbolic variational inequalities for displacements with memory-dependent constitutive laws, as seen in quasi-static viscoelasticity with self-contact [28, 42]. The latter introduces novel solution concepts distinct from classical frameworks [4, 30], emphasising admissible forces that preclude terminal-phase collisions, a paradigm central to our work as well.
@@
Recent studies highlight unresolved questions in viscoelastic obstacle problems, such as the interplay between memory kernel decay rates and solution regularity [12]. These align with broader inquiries into non-local integro-differential equations governing viscoelastic dynamics, where semigroup methods and spectral analysis remain pivotal [12]. Meanwhile, confinement conditions in linearized elasticity, particularly for three-dimensional bodies in half-spaces, demand sophisticated Sobolev-space analyses to ensure that suitable energy estimates hold [15].
In this paper, we formulate a three-dimensional Signorini problem for a time-dependent linearly viscoelastic body constrained to a half-space. To the best of our knowledge, the current literature does not address the study of time-dependent obstacle problems for three-dimensional viscoelastic bodies subjected to a confinement condition Signorini-type. Our approach synthesizes ideas from [42] - where forces appear to avoid terminal-phase contact - and the geometric confinement frameworks of [11]. The main result of this paper, namely Theorem 4.4, establishes existence of solutions under “admissible” applied body forces. The concept of solution for the governing model is recovered by two layer of approximations: in the first layer, we relax the constraint by penalising the classical energy in linearised viscoelasticity, while in the second layer we discretise the relaxed problem by Galerkin method. A noticeable feature that we establish is that the displacement acceleration enjoys the regularity of a vector-valued measure in the sense of Dinculeanu [13] or, equivalently, the displacement velocity is a function of bounded variation [29].
@@
The paper is divided into four sections, including this one. In section 2 we present the main notation from elasticity and differential geometry. In section 3 we formulate the Signorini problem for a three-dimensional viscoelastic body. Finally, in section 4, we recover the existence of solutions for the model under consideration by means of compactness methods.
@@ 2. Geometrical preliminaries
For comprehensive information on the classical concepts of differential geometry referenced in this section, see, e.g., [8] or [9].
Latin indices, except when utilized for indexing sequences, take values from the set ${1,2,3}$, and the summation convention for repeated indices is consistently applied alongside this rule.
For an open subset $\Omega\subset\mathbb{R}^{3}$, the notations $L^{2}(\Omega)$ and $H^{1}(\Omega)$ refer to the standard Lebesgue and Sobolev spaces, respectively. The notation $\mathcal{D}(\Omega)$ indicates the space of functions that are infinitely differentiable on $\Omega$ and have compact support within $\Omega$. The notation $|\cdot|_{X}$ denotes the norm in a vector space $X$. Spaces of vector-valued functions are represented by boldface letters. Lebesgue-Bochner spaces (see, e.g., [29]) are represented by the notation $L^{p}(0,T;X)$, where $1\leq p\leq\infty$, $T>0$ and $X$ is a Banach space satisfying the Radon-Nikodym property. The notation $X^{\ast}$ refers to the dual space of a vector space $X$, and the notation $\langle\cdot,\cdot\rangle_{X^{\ast},X}$ denotes the duality pairing between $X^{\ast}$ and $X$. The notation $\mathcal{C}^{0}([0,T];X)$ designates the space of continuous $X$-valued functions defined over the compact interval $[0,T]$, while the special notation $\langle\langle\cdot,\cdot\rangle\rangle_{X^{\ast},X}$ denotes the duality pair between $(\mathcal{C}^{0}([0,T];X))^{\ast}$ and $\mathcal{C}^{0}([0,T];X)$. The notation $\mathcal{M}([0,T];X^{\ast})$ denotes the space of $X^{\ast}$-valued measures defined over the compact interval $[0,T]$.
[…]
A Lipschitz domain $\Omega\subset\mathbb{R}^{3}$ is a non-empty, open, bounded, and connected subset with a Lipschitz continuous boundary $\Gamma$, where the set $\Omega$ is locally on the same side of $\Gamma$. The notation $\,\mathrm{d}x$ indicates the volume element in $\Omega$, and the symbol $\,\mathrm{d}\Gamma$ indicates the area element along $\Gamma$. Finally, let $\Gamma=\Gamma_{0}\cup\Gamma_{1}$ be a $\,\mathrm{d}\Gamma$-measurable portion of the boundary such that $\Gamma_{0}\cap\Gamma_{1}=\emptyset$ and area $\textup{area }\Gamma_{0}>0$. We denote by $\nu$ the outer-unit normal vector field along the boundary $\Gamma$ of the set $\Omega$.
As a model of the three-dimensional “physical” space $\mathbb{R}^{3}$, we consider a real three-dimensional affine Euclidean space, which is defined by selecting a point $O$ as the origin and associating it with a real three-dimensional Euclidean space, denoted $\mathbb{E}^{3}$. We equip $\mathbb{E}^{3}$ with an orthonormal basis consisting of three vectors $\bm{e}^{i}$. The Euclidean inner product of two elements $\bm{a}$ and $\bm{b}$ in $\mathbb{E}^{3}$ is denoted by $\bm{a}\cdot\bm{b}$; the Euclidean norm of any $\bm{a}\in\mathbb{E}^{3}$ is denoted by $|\bm{a}|$; the Kronecker symbol is denoted by $\delta^{ij}$.
@@
The set $\overline{\Omega}$ is the reference configuration occupied by a linearly elastic body in the absence of applied body forces. We assume that $\overline{\Omega}$ is in a natural state, meaning the body is stress-free in this configuration. Following [7], we also assume that the material is isotropic, homogeneous, and linearly elastic. Under these conditions, the behaviour of the linearly elastic material is governed by its two Lamé constants $\lambda\geq 0$ and $\mu>0$. We denote the viscosity constants by $\theta\geq 0$ and $\xi>0$. The positive constant $\rho$ designates the mass density of the linearly viscoelastic body per unit volume.
In what follows, “a.e.” stands for “almost everywhere” and “a.a.” stands for “almost all”. Define the space
$\bm{V}(\Omega):=\{\bm{v}=(v_{i})\in\bm{H}^{1}(\Omega);\bm{v}={\bf{0}}\textup{ on }\Gamma_{0}\},$
@@
We define the three-dimensional elasticity tensor in Cartesian coordinates and we denote its components by $A^{ijk\ell}$. We recall that the contravariant components of this tensor are defined by (see, e.g., [7])
@@
We then define the three-dimensional viscoelasticity tensor in Cartesian coordinates and we denote its components by $B^{ijk\ell}$. We recall that the contravariant components of this tensor are defined by (see, e.g., [22])
@@
For each $\bm{v}\in\bm{H}^{1}(\Omega)$ we consider the linearised change of metric tensor $\bm{e}(\bm{v})$, whose components $e_{i\|j}(\bm{v})$ are defined by:
$e_{i\|j}(\bm{v}):=\dfrac{1}{2}(\partial_{j}v_{i}+\partial_{i}v_{j})\in L^{2}(\Omega).$
This tensor is symmetric, i.e., $e_{i\|j}(\bm{v})=e_{j\|i}(\bm{v})$, for all $\bm{v}\in\bm{H}^{1}(\Omega)$.
We now state Korn’s inequality in Cartesian coordinates (see, e.g., Theorem 6.3-6 of [7]).
@@
Let $\Omega$ be a Lipschitz domain in $\mathbb{R}^{3}$ and let $\Gamma_{0}$ be a non-zero area subset of the whole boundary $\Gamma$. Then, there exists a constant $c_{0}>0$ such that
$c_{0}^{-1}\|\bm{v}\|_{\bm{H}^{1}(\Omega)}\leq\|\bm{e}(\bm{v})\|_{\bm{L}^{2}(\Omega)}\leq c_{0}\|\bm{v}\|_{\bm{H}^{1}(\Omega)},$
@@
We denote by $\mathcal{A}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ the linear and continuous operator defined in a way that:
$\langle\mathcal{A}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,\quad\textup{ for all }\bm{w},\bm{v}\in\bm{V}(\Omega).$
Thanks to the properties of the fourth order three-dimensional elasticity tensor $\{A^{ijk\ell}\}$ and Korn’s inequality in Cartesian coordinates (Theorem 2.1), we have that the operator $\mathcal{A}$ is symmetric and uniformly positive-definite, in the sense that there exists a constant $C_{e}^{(1)}>0$ such that:
@@
In a completely similar fashion, we denote by $\mathcal{B}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ the linear and continuous operator defined in a way that:
$\langle\mathcal{B}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,\quad\textup{ for all }\bm{w},\bm{v}\in\bm{V}(\Omega).$
Thanks to the properties of the fourth order three-dimensional elasticity tensor $\{B^{ijk\ell}\}$ and Korn’s inequality in Cartesian coordinates (Theorem 2.1), we have that the operator $\mathcal{B}$ is symmetric and uniformly positive-definite (cf., e.g., [22]), in the sense that there exists a constant $C_{e}^{(2)}>0$ such that:
@@ 3. Variational formulation of the time-dependent obstacle problem for a linearly viscoelastic body
In this paper, we examine a specific obstacle problem for a linearly elastic body subjected to a confinement condition. This condition requires that any admissible displacement vector field $v_{i}\bm{e}^{i}$ must ensure that all points of the corresponding deformed configuration remain within a half-space of the form
:=$\mathbb{H}:=\{x\in\mathbb{R}^{3};\bm{Ox}\cdot\bm{q}\geq 0\},$
where $\bm{q}$ is a unit-norm vector given once and for all. Let us denote by $\bm{I}$ the identity mapping $\bm{I}:\overline{\Omega}\to\mathbb{E}^{3}$ and let us assume that the undeformed reference configuration satisfies
$\bm{I}(x)\cdot\bm{q}\geq 0,\quad\textup{ for all }x\in\overline{\Omega},$
or, in other words, there may be contact between the obstacle and the reference configuration when no applied forces are acting on the reference configuration.
@@
for all $x\in\Gamma$ or, possibly, only for a.a. $x\in\Gamma$ when the covariant components $v_{i}$ are required to belong to the space $H^{1}(\Omega)$. The latter is a confinement condition of Signorini type (cf., e.g., [27]).
Let us introduce the subset $\bm{U}(\Omega)$ of admissible displacements takes the following form:
$\bm{U}(\Omega):=\{\bm{v}=(v^{i})\in\bm{V}(\Omega);(\bm{I}+v_{i}\bm{e}^{i})\cdot\bm{q}\geq 0\,\textup{a.e. on }\Gamma\}.$
The linearly viscoelastic body under consideration is subjected to both applied body forces, that we denote by $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$. The applied surface forces for the problem under consideration are expressed in terms of the vectors $\bm{e}^{i}$ of the Cartesian framework. For the sake of simplicity, we do not consider the action of surface forces, even though it is licit to consider the action of traction forces on the portion of the boundary that does not engage contact with the obstacle.
Moreover, we restrict ourselves to considering applied body forces in the sense of the definition stated below.
Definition 3.1.
We say that an applied body force $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$ is admissible if there exists $T_{0}=T_{0}(\bm{f})\in(0,T)$ such that the constraint becomes inactive for a.a. $T_{0}<t<T$.
By restricting ourselves to considering applied body forces that are admissible in the sense of Definition 3.1, we are allowed to incorporate this information in the governing model for the problem under consideration, and we thus solely need to focus on making sure that the constraint is implemented for a.a. $0<t<T_{0}$. This specialisation of the applied body forces will be critical for establishing the sought existence result. In connection with this assumption, we mention the recent paper [42] where the authors considered a variational formulation of a collision problem non-linearly viscoelastic solids. The concept of solution proposed in [42] suggests that the applied force shuts down and that the terminal point $T$ is so large so as to have zero velocity in terminal part of the motion. The case considered in [42] is a particular case of “admissible” applied body force in the sense of Definition 3.1.
In the problem here discussed, we propose a different concept of solution that is more similar to the classical concept of solutions for variational inequalities [6, 19, 30]. In our formulation, the velocity of the linearly viscoelastic body needs not vanishing in the long run; however, we need to incorporate in the governing model the requirement that the constraint is inactive in the terminal part of the motion.
@@
In all what follows, we restrict ourselves to considering applied body forces that are “admissible” in the sense of Definition 3.1.
@@
$\dot{\bm{u}}\in L^{\infty}(0,T;\bm{L}^{2}(\Omega))\cap L^{2}(0,T;\bm{V}(\Omega)),$
$\ddot{\bm{u}}\in\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega)),$
that satisfies the following variational inequalities
$2\rho\langle\langle\ddot{\bm{u}},\bm{v}\rangle\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}-2\rho\langle\dot{\bm{u}}(T),\bm{u}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+2\rho\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+2\rho\|\dot{\bm{u}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}$
@@
$\quad-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}(T))e_{i\|j}(\bm{u}(T))\,\mathrm{d}x\,\mathrm{d}t+\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x$
@@
for all $\bm{v}=(v_{i})\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$, and that satisfies the following initial conditions
@@
where $\bm{u}_{0}=(u_{i,0})\in\bm{U}(\Omega)$, and $\bm{u}_{1}=(u_{i,1})\in\bm{V}(\Omega)$ are prescribed. $\blacksquare$

### v3

# December 2, 2025

@@ Abstract.
In this paper we establish the existence of solutions for a model describing the evolution of a linearly viscoelastic body which is constrained to remain confined in a prescribed half-space. The confinement condition under consideration is of Signorini type, and is given over the boundary of the linearly viscoelastic body under consideration. We show that one such variational problem admits solutions and we coin a novel concept of solution which, differently from the available literature, is valid even in the case where the viscoelastic body starts its motion in contact with the obstacle. Additionally, under additional assumptions on the constituting material, we show that when the applied body force is lifted the deformed linearly viscoelastic body returns to its rest position at an exponential rate of decay.
@@ 1. Introduction
The contact of elastic and viscoelastic materials with rigid obstacles arises in many applicative fields, including biomedical engineering, aerospace design, and materials science. For instance, the motion of aortic heart valves - modelled as linearly elastic shells constrained within a spatial domain to avoid collisions - has inspired mathematical frameworks for obstacle problems in elasticity [29, 33, 50, 53]. Similarly, the confinement of deformable structures like polymer films or viscoelastic membranes in prescribed half-spaces is critical for applications ranging from soft robotics to biomechanical implants [49]. Recently, the theory of obstacle problems was used to model ice melting sheets [23, 34, 46].
@@
The mechanical deformation of linearly viscoelastic bodies is classically modelled using energy functionals from linearized elasticity theory [12, 29]. Recent advances extend these principles to dynamic and time-dependent settings. For example, Bock and Jarušek analysed unilateral dynamic contact problems for viscoelastic Reissner-Mindlin plates [5], while [7] established existence results for thermoelastic von Kármán plates vibrating against rigid obstacles. A key challenge in such problems lies in reconciling hyperbolic variational inequalities for displacements with memory-dependent constitutive laws, as seen in quasi-static viscoelasticity with self-contact [36, 55]. The latter introduces novel solution concepts distinct from classical frameworks [8, 38], emphasising admissible forces that preclude terminal-phase collisions.
@@
Recent studies highlight unresolved questions in viscoelastic obstacle problems, such as the interplay between memory kernel decay rates and solution regularity [18]. These align with broader inquiries into non-local integro-differential equations governing viscoelastic dynamics, where semigroup methods and spectral analysis remain pivotal [18]. Meanwhile, confinement conditions in linearized elasticity, particularly for three-dimensional bodies in half-spaces, demand sophisticated Sobolev-space analyses [20].
For what concerns the study of the dynamics linearly elastic bodies subjected to remaining confined in a prescribed half-space, we refer to [45]. The existence result in this paper is established by resorting to vector-valued measures in the sense of Dinculeanu. The latter approach requires some strong assumptions, which we aim to overcome in the present paper.
In this paper, we formulate a three-dimensional obstacle problem for a time-dependent linearly viscoelastic body constrained to a half-space. To the best of our knowledge, the current literature does not address the study of time-dependent obstacle problems for three-dimensional viscoelastic bodies subjected to a confinement condition Signorini-type. For completeness, we mention the textbook by Han & Sofonea [29] where quasi-static viscoelastic problems are studied by means of a different concept of solution from the one employed here. Additionally, we mention the recent papers [2, 3, 58] where the quasi-static Signorini problems in viscoelasticity are considered from the analytical and numerical point of view.
Our approach synthesizes ideas from [55], where forces appear to avoid terminal-phase contact, and the geometric confinement frameworks of [17]. Critical to establish the existence of solutions for the problem considered here is the compactness for the time-dependent version of the trace operator (Lemma 4.2), which appears to be a novel result to the best of our knowledge. We also mention the recent paper [56], where the authors consider a Signorini problem in linearised viscoelasticity with a contact condition for both the displacement and the velocity. In this contribution, the authors penalise the problem and obtain estimates depending on a parameter.
In addition to the latter, the novelties introduced in this paper are:
  • $(1)$
    The derivation of a sound concept of solution for a time-dependent obstacle problem of Signorini type in linearised viscoelasticity, which aligns to the classical concept of solution for obstacle problems for elliptic models;
  • $(2)$
    The methodology here presented, differently from [55], takes an initial condition for the velocity into consideration. Additionally, our proposed methodology does not take into account the terminal condition implicitly assumed in [55];
  • $(3)$
    Our methodology takes into account a confinement condition of Signorini type rather than the Ciarlet-Nečas condition (see [16]) considered in [55];
  • $(4)$
    Differently from [4, 5, 6, 7], the methodology here presented applies to obstacle problems formulated in terms of vectorial magnitudes;
  • $(5)$
    By contrast with [56], we only impose the constraint on the displacement and we consider non-zero initial data and we do not assume additional regularity for the solution;
  • $(6)$
    By contrast with the recent works [17, 45], our methodology has the advantage of applying to bodies which start their motion in contact with the obstacle. Moreover, all the extra assumptions made in [45] are dropped, suggesting that viscoelastic models are more suitable for studying the dynamics of elastic bodies making contact with obstacles during their motion. This statement is corroborated by the conclusion derived in Theorem 4.5, where we show that when the applied body force is released the body tends to return to its original reference configuration.
The concept of solution for the governing model is recovered by two layer of approximations: in the first layer, we relax the constraint by penalising the classical energy in linearised viscoelasticity, while in the second layer we discretise the relaxed problem by Galerkin method.
@@
The paper is divided into four sections, including this one. In section 2 we present the main notation from elasticity and differential geometry. In section 3 we formulate the obstacle problem for a three-dimensional viscoelastic body. In section 4, we recover the existence of solutions for the model under consideration by means of compactness methods.
@@ 2. Geometrical preliminaries
For an open subset $\subset\mathbb{R}^{3}$, the notations $L^{2}(\Omega)$ and $H^{1}(\Omega)$ refer to the standard Lebesgue and Sobolev spaces, respectively. The notation $\mathcal{D}(\Omega)$ indicates the space of functions that are infinitely differentiable on $\Omega$ and have compact support within $\Omega$. The notation $|\cdot|_{X}$ denotes the norm in a vector space $X$. Spaces of vector-valued functions are represented by boldface letters like for instance $\bm{L}^{2}(\Omega)$ or $\bm{H}^{1}(\Omega)$. blueSpaces of symmetric tensors are denoted in blackboard bold letters like for instance .
Lebesgue-Bochner spaces (see, e.g., [37]) are represented by the notation , where $1\leq p\leq\infty$, $T>0$ and $X$ is a Banach space satisfying the Radon-Nikodym property. The notation refers to the dual space of a vector space $X$, and the notation denotes the duality pairing between $X^{\ast}$ and $X$. The notation designates the space of continuous $X$-valued functions defined over the compact interval $[0,T]$.
The notations and represent the first weak derivative with respect to $t\in(0,T)$ and the second weak derivative with respect to $t\in(0,T)$ of a scalar function $\eta$ defined over the interval $(0,T)$. The notations and denote the first weak derivative with respect to $t\in(0,T)$ and the second weak derivative with respect to $t\in(0,T)$ of a vector field $\bm{\eta}$ defined over the interval $(0,T)$.
A Lipschitz domain $\Omega\subset\mathbb{R}^{3}$ is a non-empty, open, bounded, and connected subset with a Lipschitz continuous boundary , where the set $\Omega$ is locally on the same side of $\Gamma$. The notation indicates the volume element in $\Omega$, and the symbol indicates the area element along $\Gamma$. For more details about this definition see, for instance, Section 8.2 in [15]. Finally, let $\Gamma={{}}{}\cup{{}}{}$ be a $\mathrm{d}\Gamma-$measurable portion of the boundary such that $\Gamma_{0}\cap\Gamma_{1}=\emptyset$ and $\textup{area }\Gamma_{0}>0$.
As a model of the three-dimensional “physical space” $\mathbb{R}^{3}$, we consider a real three-dimensional affine Euclidean space, which is defined by selecting a point $O$ as the origin and associating it with a real three-dimensional Euclidean space, denoted . We equip $\mathbb{E}^{3}$ with an orthonormal basis consisting of three vectors . The Euclidean inner product of two elements $\bm{a}$ and $\bm{b}$ in $\mathbb{E}^{3}$ is denoted by ; the Euclidean norm of any $\bm{a}\in\mathbb{E}^{3}$ is denoted by ; the Kronecker symbol is denoted by .
@@
The set $\overline{\Omega}$ is the reference configuration occupied by a linearly viscoelastic body in the absence of applied forces. We assume that $\overline{\Omega}$ is in a natural state, meaning the body is stress-free in this configuration. Following [12], we also assume that the material is isotropic, homogeneous, and linearly viscoelastic. Under these conditions, the behaviour of the linearly elastic material is governed by its two Lamé constants $\geq 0$ and >0{{}}{}>0. We denote the viscosity constants by $\geq 0$ and >0{{}}{}>0. The positive constant designates the mass density of the linearly viscoelastic body per unit volume.
In what follows, “” stands for “almost everywhere” and “” stands for “almost all”. Define the space
$\{\bm{v}=(v_{i})\in\bm{H}^{1}(\Omega);\bm{v}={\bf{0}}\,\,\mathrm{d}\Gamma\textup{-a.e. on }\Gamma_{0}\},$
@@
Note that in the definition of the space $\bm{V}(\Omega)$ we require the trace of any element $\bm{v}$ to vanish on $\Gamma_{0}$ up to a zero-measure subset of $\Gamma_{0}$ with respect to the measure of surfaces. The trace operator is denoted and defined by $\bm{V}(\Omega)\to\bm{L}^{2}(\Gamma)$, and we recall that tr is linear, continuous and compact [40].
Define the fourth order three-dimensional elasticity tensor in Cartesian coordinates and denote its components by . We recall that the contravariant components of this tensor are defined by (see, e.g., [12])
@@
We then define the fourth order three-dimensional viscosity tensor in Cartesian coordinates and we denote its components by . We recall that the contravariant components of this tensor are defined by (see, e.g., [29])
@@
For each $\bm{v}\in\bm{H}^{1}(\Omega)$ we consider the linearised change of metric tensor in Cartesian coordinates $\bm{e}(\bm{v})$, whose components $e_{i\|j}(\bm{v})$ are defined by:
$(\bm{v}):=\dfrac{1}{2}(\partial_{j}v_{i}+\partial_{i}v_{j})\in L^{2}(\Omega).$
This tensor is symmetric, i.e., $e_{i\|j}(\bm{v})=e_{j\|i}(\bm{v})$, for all $\bm{v}\in\bm{H}^{1}(\Omega)$. We now state Korn’s inequality in Cartesian coordinates (see, e.g., Theorem 6.3-4 of [12]).
@@
Let $\Omega$ be a Lipschitz domain in $\mathbb{R}^{3}$ and let $\Gamma_{0}$ be a non-zero area subset of the whole boundary $\Gamma$. Then, there exists a constant >0{{}}{}>0 such that
$c_{0}^{-1}\|\bm{v}\|_{\bm{H}^{1}(\Omega)}\leq\|\bm{e}(\bm{v})\|_{\mathbb{L}^{2}(\Omega)}\leq c_{0}\|\bm{v}\|_{\bm{H}^{1}(\Omega)},$
@@
We denote by $\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ the linear and continuous operator defined in a way that:
(1) $\langle\mathcal{A}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,\quad\textup{ for all }\bm{w},\bm{v}\in\bm{V}(\Omega).$
Thanks to the properties of the fourth order three-dimensional elasticity tensor $\{A^{ijk\ell}\}$ it results that the operator $\mathcal{A}$ is symmetric and uniformly positive-definite, in the sense that there exists a constant >0{{}}{}>0 such that:
@@
In a completely similar fashion, we denote by $\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ the linear and continuous operator defined in a way that:
(2) $\langle\mathcal{B}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,\quad\textup{ for all }\bm{w},\bm{v}\in\bm{V}(\Omega).$
Thanks to the properties of the fourth order three-dimensional viscosity tensor $\{B^{ijk\ell}\}$ it results that the operator $\mathcal{B}$ is symmetric and uniformly positive-definite (cf., e.g., [29]), in the sense that there exists a constant >0{{}}{}>0 such that:
@@ 3. Variational formulation of the time-dependent obstacle problem for a linearly viscoelastic body
In this section, we formulate a specific obstacle problem for a linearly viscoelastic body subjected to remaining confined in a prescribed half-space. This condition requires that any admissible displacement vector field $v_{i}\bm{e}^{i}$ must ensure that all points of the corresponding deformed configuration remain within a half-space of the form
$\{x\in\mathbb{R}^{3};\bm{Ox}\cdot\bm{q}\geq 0\},$
where is a unit-norm vector that is given once and for all. Let us denote by the identity mapping $\bm{I}:\overline{\Omega}\to\mathbb{E}^{3}$ and let us assume that the undeformed reference configuration satisfies
(3) $\bm{I}(x)\cdot\bm{q}\geq 0,\quad\textup{ for all }x\in\overline{\Omega},$
or, in other words, the linearly viscoelastic body is located in the prescribed half-space when at rest and contact may occur. Note that (3) does not imply that the body does not start its motion in contact with the obstacle. The latter property is, in fact, associated with the initial condition.
@@
for all $x\in\Gamma$ or, possibly, only for a.a. $x\in\Gamma$ when the covariant components $v_{i}$ are required to belong to the space $H^{1}(\Omega)$. The latter is a confinement condition of Signorini type (cf., e.g., [35]) and means that the boundary of the linearly viscoelastic body under consideration cannot cross the prescribed half-space identified by its orthonormal complement $\bm{q}$. This condition is widely accepted to imply that all the other points in the deformed reference configuration of the linearly elastic body under consideration must not cross the prescribed half-space.
The subset of admissible displacements is defined by:
$\bm{U}(\Omega):=\{\bm{v}=(v^{i})\in\bm{V}(\Omega);(\bm{I}+v_{i}\bm{e}^{i})\cdot\bm{q}\geq 0\,\,\mathrm{d}\Gamma\textup{-a.e. on }\Gamma\}.$
The linearly viscoelastic body under consideration is subjected to applied body forces, that we denote by $(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$. The applied surface forces for the problem under consideration are expressed in terms of the vectors $\bm{e}^{i}$ of the Cartesian framework. For the sake of simplicity, we do not consider the action of tractions, even though it is licit to consider the action of traction forces on the portion of the boundary that does not engage contact with the obstacle. Further in the analysis, we will make additional assumptions on the applied body force $\bm{f}$, requiring that $\bm{f}\in H^{1}(0,T;\bm{L}^{2}(\Omega))$.
In the problem here discussed, we propose a different concept of solution that is more similar to the classical concept of solutions for variational inequalities [9, 26, 38]. In our formulation, the velocity of the linearly viscoelastic body needs not vanishing in the long run, as it seems to be postulated in [55].
@@
$\dot{\bm{u}}\in\mathcal{C}^{0}([0,T];\bm{L}^{2}(\Omega))\cap L^{2}(0,T;\bm{V}(\Omega)),$
that satisfies the variational inequalities
$2\rho\int_{\Omega}\dot{\bm{u}}(T)\cdot(\bm{v}(T)-\bm{u}(T))\,\mathrm{d}x-2\rho\int_{\Omega}\bm{u}_{1}\cdot(\bm{v}(0)-\bm{u}_{0})\,\mathrm{d}x-2\rho\int_{0}^{T}\int_{\Omega}\dot{\bm{u}}(t)\cdot(\dot{\bm{v}}(t)-\dot{\bm{u}}(t))\,\mathrm{d}x\,\mathrm{d}t$
@@
$\quad-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}(T))e_{i\|j}(\bm{u}(T))\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x$
@@
for all $\bm{v}=(v_{i})\in L^{\infty}(0,T;\bm{U}(\Omega))$ such that $\dot{\bm{v}}\in L^{\infty}(0,T;\bm{L}^{2}(\Omega))\cap L^{2}(0,T;\bm{V}(\Omega))$, and that satisfies the initial conditions
@@
where $(u_{i,0})\in\bm{U}(\Omega)$, and $(u_{i,1})\in\bm{L}^{2}(\Omega)$ are prescribed. $\blacksquare$
Let us note in passing that the initial condition $\bm{u}_{0}$ is required to belong to the set $\bm{U}(\Omega)$. This means that the linearly viscoelastic body under consideration could make contact with the obstacle at the beginning of the observation. This feature could not be considered in [45, 47] due to the intrinsic limitations of the model.